Pleas help solve the equation.
8 + 4 ln x = 5

Answer by stanbon(57337)

Pleas help solve the equation.
8 + 4 ln x = 5
----
4*ln(x) = -3
ln(x) = -3/4
----
x = e^(-3/4)
x = 0.4724
===================
cheers,
Stan H.
===================